Circular No. 6402

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

1996 JA_{1}

MPEC 1996-K01 gives word of the discovery, by T. Spahr at the University of Arizona on May 14.4 UT, of a bright asteroidal object rapidly approaching the earth. The extensive follow-up observations (published also on MPEC 1996-K03 and -K04) show that the object (a = 2.53 AU, e = 0.70, i = 22^{o}) will pass only 0.00303 AU from the center of the earth on May 19.690 UT. With absolute visual magnitude H = 21.0, 1996 JA_{1} is probably the largest such object ever observed to approach the earth so closely. The object should brighten to V ∼ 11, and it will be necessary for observers to allow for parallax in their ephemerides; a facility for computing topocentric ephemerides at UT intervals as short as 10 min is available in

the WWW at http://cfa-www.harvard.edu/cfa/ps/NEO/1996JA1.html.

D. J. Tholen, Institute of Astronomy, University of Hawaii, reports that his single-filter photometry with the 2.2-m reflector showed 1996 JA_{1} to brighten by ∼ 0.75 mag between May 17.254 UT and what was a fairly broad maximum around May 17.311. The object was almost 0.5 mag fainter by May 17.353, indicating that the rotation period was rather more than 5.5 hr. On the following night, he observed a broad maximum around May 18.297 and a sharp minimum on May 18.338, suggesting eight cycles between the maxima, i.e., a rotation period of 5.92 hr. By May 18.353 there had again been brightening by 0.3 mag. The 0.75-mag range between the May 18 maximum and minimum suggests that the object was very close to minimum when first observed on May 17.254; nine cycles between then and May 18.338 correspond to a rotation period of 5.78 hr, in plausible agreement with the value established from the maxima. Calibration using a set of solar-color Landolt standards confirmed that the adopted H = 21.0 (with slope parameter G = 0.15) is close to the object’s mean value.

M. Hicks and J. Collins, Lunar and Planetary Laboratory, University of Arizona, report: “CCD spectra (range 0.55–1.05 μm, resolution 1.5 nm) were obtained with the Catalina Station 1.54-m telescope on May 17 and 18. Preliminary reductions of the first night’s data indicate that 1996 JA_{1} is of Tholen’s V type (Tholen and Barucci 1989, Asteroids II, p. 298; Binzel and Xu 1993, Science 260, 186). The relative depth of the pyroxene absorption band centered at 0.92–0.94 μm is 0.75 for 1996 JA_{1}, compared to 0.72 for (4) Vesta. The spectrum also shows the characteristic reddish slope, increasing by 18 percent over the range 0.55–0.70 μm.”

1996 May 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden